A spaceship is moving away from Earth with speed 0.03c where c is speed of light in air.The spaceship emits signal of frequency 6×10^(7)Hz. Frequency of signal observed by the observer on Earth will be
1
Expert's answer
2021-09-08T16:13:50-0400

Given:

"v=-0.03c"

"f_0=6*10^7\:\rm Hz"


The Doppler shift of a frequency

"f=f_0\sqrt{\frac{1+v/c}{1-v/c}}"

So

"f=6.00*10^7\sqrt{\frac{1-0.03}{1+0.03}}=5.82*10^7\:\rm Hz"